OpenStudy (anonymous):

what is the quotient of 2/3 divided by 3/10 in simplest form? 1/5 5/13 2 2/9 3 1/3

OpenStudy (anonymous):

(2/3)/(3/10)=(2/3)*(10/3)=20/9 20/9 is 2 with 2 left over... so 2 2/9

OpenStudy (anonymous):

thanks. :)